# Chatbots and Zero Sales Resistance

### v1

@@ 3 Criticism to ML
There are indeed several solid reasons to second-guess the blue-sky scenarios portrayed by the most enthusiastic (and aggressive ) AI aficionados. Here we briefly summarize the main items, some of which have been pointed out before [11, 12]:
@@ 4 Sustainability
Besides the exponential number of parameters, in order to converge, the ML procedure needs correspondingly huge training datasets. As mentioned above, training does not come for free: it is estimated that new generation chatbots will come near to the Gigawatt power demand, in excess of most existing power plants. The comparison with the 20 Watts of our brain is embarrassing. True, once trained, their routine operation is much less expensive. The question still remains, as to whether, chatbots can still function accurately and reliably on unseen data.
@@
This is the biggest lie which has been served to us by AAI, where AAI stands for Aggressive AI, in order to distinguish it from the many important contributions of AI to technology and society. But the king is now naked, for the pursuit of Control regardless of Insight comes with a energy price tag that planet Earth just cannot sustain. Incidentally, scientists (most physicists and mathematicians) much less visible than Musk and LeCun, are already doing this, out of the limelight [16, 17].
@@ 7 Summary
Machine learning has potential to provide a substantial new entry to the world of scientific investigation, as it can occasionally capture levels of complexity beyond reach for our best theoretical and computational models. However the blind pursuit of ever-increasing number of weights with little or no attention to Insight versus Control, is overly dangerous in several respects. Besides being energetically unsustainable, it paves the way to zero-sales-resistance strategies whereby Science is served as a perfect strawman for Control and financial power.

### v2

@@ 3 Criticism to ML
There are indeed several solid reasons to second-guess the blue-sky scenarios portrayed by the most enthusisatic (and aggressive ) AI aficionados. Here we briefly summarize the main items, some of which have been pointed out before [11, 12]:
@@ 4 Sustainability
Besides the exponential number of parameters, in order to converge, the ML procedure needs correspondingly huge training datasets. As mentioned above, training does not come for free: it is estimated that new generation chatbots will come near to the Gigawatt power demand, in excess of most existing power plants. The comparison with the 20 Watts of our brain is embarassing. True, once trained, their routine operation is much less expensive. The question still remains, as to whether, chatbots can still function accurately and reliably on unseen data.
@@
This is the biggest lie which has been served to us by AAI, where AAI stands for Aggressive AI, in order to distinguish it from the many important contributions of AI to technology and society. But the king is now naked, for the pursuit of Control regardless of Insight comes with a energy price tag that planet Earth just cannot sustain. Incidentally, scientists (most physicists and mathematicians) much less visible than Musk and LeCun, are already doing this, out of the limelight [16, 17, 18].
@@ 7 Summary
Machine learning has potential to provide a substantial new entry to the world of scientific investigation, as it can occasionally capture levels of complexity beyond reach for our best theoretical and computational models. However the blind pursuit of ever-increasing number of weights with little or no attention to Insight versus Control, is overly dangerous in several respects. Besides being energetically unsustainable, it paves the way to zero-sales-resistence strategies whereby Science is served as a perfect strawman for Control and financial power.